a parallelogram is formed with a and b as the sides.
let d1 and d2 be the diagonals of the parallelogram, then a^2 + b^2 =?
(in terms of vector product)

Grade:11

2 Answers

Vikas TU
4 years ago
Hiii Numa
Parallelogram Law:
If two vector acting on a particle at the same time are represented in magnitude and directed in two adjacent sides od paralleloram drawn from a pint  this resultant vector are represented by.:
a^2 + b^2 = (a+b) (a+b) = d1*d2
aswanth nayak
4 years ago
Dear Student,
 Here ,According to Parallelogram law ,
Parallelogram Law:
If two vector acting on a particle at the same time are represented in magnitude and directed in two adjacent sides od paralleloram drawn from a pint  this resultant vector are represented by.
a^2 + b^2 = (a+b) (a+b) = d1*d2
i.e.;a^2=(a+b),b^2=(a+b)
Hope it helps you
regards